1.The position vector of a particle is given as r?=(t2-4t+6)î+(t2)j.The time after which the velocity vector and accleration vector becomes perpendicular to each other is equal to

Grade:12

3 Answers

Ashutosh Gupta
12 years ago

never

 

KARAN ACHARYA
12 years ago

v = dr/dt = (2t-4)i +(2t)j       v = velocity

a = 2i+2j                          a = acceleration

 

for a and v to be perpendicular,

=>  a.v = 0           . is dot product

=>   (2i+2j).[(2t-4)i+(2t)j] = 0

=>     4t -8 +4t = 0

=>     8t =8 

i.e t=1 

Krishna Ujjwal
6 years ago
A block of mass M is moving on a smooth horizontal surface with constant speed u. Bullets arefired horizontally against the block to reduce the velocity of the block to half its initial value.Bullets get embedded in the block. Mass of each bullet is ‘m’ and speed ‘u’. Then the number ofbullets required is–